# https://semantic-gaussians.github.io

### v1

@@ 4 Experiments
In this section, we conduct various experiments to demonstrate the effectiveness of Semantic Gaussians on open-vocabulary 3D scene understanding and other applications. We first evaluate and compare our method on the ScanNet 2D semantic segmentation benchmark which is constructed on 3D scenes. Then, we exhibit qualitative results on some applications, including part segmentation, spatiotemporal tracking, and language-guided editing. We also provide ablation study results.
4.1 Experimental Setup
@@
For comparisons on semantic segmentation task, we select ScanNet [6] dataset as our 2D segmentation benchmark. ScanNet is a large-scale segmentation benchmark of indoor scenes, with calibrated RGBD trajectories, 3D point clouds, and ground-truth semantic label maps. We train 3D RGB Gaussians for all 1,201 scenes. We train our 3D semantic network on ScanNet train set and evaluate our performance on the validation set.
@@
For qualitative evaluations, we choose MVImgNet [35] dataset as our part segmentation dataset, and CMU Panoptic [13] dataset as our spatiotemporal tracking dataset. MVImgNet is a multi-view single-object dataset containing 238 classes of objects with camera parameters and sparse point clouds. CMU Panoptic dataset is a large-scale dataset for multi-people engaging in social activities. To evaluate Semantic Gaussians on 4D Gaussians, we follow the work of Dynamic 3D Gaussians [24] to construct Gaussians with temporal information. For language-guided editing, we choose some scenes in the Mip-NeRF 360 [2] dataset to show our performance.
4.2 Semantic Segmentation
4.2.1 Comparisons on ScanNet-20.
We first evaluate our approach on the 2D semantic segmentation task. We compare our method with OpenSeg [10], an open-vocabulary 2D segmentation model. We report mIoU and mAcc as the metrics of semantic segmentation on the ScanNet-20 benchmark. As we only process off-the-shelf 3D Gaussians and do not care about their training process, we do not report any metrics about rendering quality. The training of 3D Gaussian Splatting follows the official setting. As we are only concerned about semantic segmentation performance, in our method, we extract semantic features from OpenSeg to conduct 2D projection and 3D network training. Consequently, the result in this section will show how much our method will improve from 2D vision-language models.
@@ 4.2 Semantic Segmentation
Table 1: 2D semantic segmentation results on the ScanNet-20 validation set. We report the mean IoU and the mean accuracy on all 20 classes.
mIoU mAcc
OpenSeg 45.3 62.0
2D Projection $s^{\text{2D}}$ 48.3 65.8
3D Semantic Network $s^{\text{3D}}$ 47.2 62.9
2D+3D Ensemble 49.5 66.0
We compare our method to the ScanNet-20 segmentation benchmark. ScanNet-20 benchmark has 20 canonical classes (wall, floor, table, etc.) with ground truth semantic labels on each posed 2D image. Table 1 shows the performance of the OpenSeg 2D model and our Semantic Gaussians. We observe that both our 2D projection (Sec. 3.1) and our 3D network (Sec. 3.2) surpass the pre-trained OpenSeg model, even though all of our knowledge comes from it and relies on no ground truth labels. We assume the reason lies in the multi-view information integration, keeping the prediction consistent and mitigating some errors in low-quality views. Moreover, we notice that though the mIou and mAcc of our 3D network are lower than the 2D projection, the 2D and 3D ensemble will further improve our performance. We conjecture that some objects in certain scenes cannot be correctly recognized by 2D models in all views due to their low quality, while the 3D network could recognize them by utilizing geometric details.
4.2.2 Ablation Study.
We conduct ablation studies to figure out the efficiency of our method. As the performance of 2D projection and 3D network is presented in Sec. 4.2.1, in this section, we ablate the performance of reducing input features to our 3D network. Points in 3D Gaussian Splatting have many features, i.e., coordinates, colors, rotations, scales, and opacities. If we only use coordinates and colors, they degrade to point clouds. Table 2 shows the comparisons of different inputs. We observe that if we reduce the input features, the mIoU and mAcc of our 3D networks will become much lower, even lower than the OpenSeg pre-trained model in Table 1. This implies that the extra features in 3D Gaussian Splatting are important, and provide more information than RGB point clouds.
Table 2: Performance of different input features on the ScanNet-20 validation set. "XYZ+RGB" means the input of the 3D network only contains the coordinate and the color information of 3D Gaussians, and "Full Gaussians" keeps the same setting as our main experiments. "3D" and "2D+3D" denote our 2D projection method and the 2D-3D ensemble method.
mIoU mAcc
3D Only XYZ+RGB 44.8 59.2
Full Gaussians 47.2 62.9
2D+3D Only XYZ+RGB 47.7 63.9
Full Gaussians 49.5 66.0
4.3 Part Segmentation
@@
4.4 Spatiotemporal Tracking
(a) basketball
(b) juggle
Figure 4: Qualitative results of spatiotemporal tracking on the CMU Panoptic dataset.
(a) softball
(b) tennis
Figure 5: More qualitative results of spatiotemporal tracking on the CMU Panoptic dataset.
@@ 4.4 Spatiotemporal Tracking
Fig. 4 and Fig. 5 show the qualitative result of our spatiotemporal tracking. As the scene often contains 1 2 humans and a dynamic object, we use VLPart in 2D projection and perform human part segmentation simultaneously. For each dynamic 3D Gaussian, we conduct our method at every frame to avoid looking at future frames. We render some novel views to show our capability of spatial tracking. The results show that our Semantic Gaussians can track human parts and objects with a high accuracy between different views and timesteps.
4.5 Language-guided Editing
@@
In this section, we show the application of language-guided editing of our Semantic Gaussians. Our method can predict semantic embeddings for each 3D Gaussian point, and thus we can choose certain points by language query. Fig. 6 shows some qualitative examples of language-guided editing on the room scene in the Mip-NeRF 360 dataset [2]. In this example, we use CLIP text embeddings of "glass bottle", "metal bowl" and "floor" to query and choose certain 3D Gaussian points in the scene, and we edit them by modifying their properties such as coordinates, colors, opacities, etc. Examples in Fig. 6 demonstrate our effectiveness in language-guided editing.
5 Conclusion
In this work, we propose Semantic Gaussians, a novel approach to open-vocabulary 3D scene understanding via 3D Gaussian Splatting. Semantic Gaussians distill knowledge from pre-trained 2D encoders by projecting 2D pixel-level embeddings to 3D Gaussian points. Moreover, we introduce a 3D sparse convolutional network to predict semantic components with the input of RGB Gaussians, thus achieving zero-shot generalization to unseen 3D scenes. We conduct experiments on the ScanNet segmentation benchmark to prove its effectiveness and exhibit downstream applications such as part segmentation, spatiotemporal tracking, and scene editing. Our work paves the way for real-world applications of 3D Gaussian Splatting, such as embodied agents and augmented reality systems.

### v2

@@ IV Experiments
In this section, we conduct various experiments to demonstrate the effectiveness of Semantic Gaussians on open-vocabulary 3D scene understanding and other applications. We first evaluate and compare our method on the ScanNet 2D semantic segmentation benchmark which is constructed on 3D scenes. Then, we exhibit the 3D object localization results of Semantic Gaussians on LERF dataset. Next, we exhibit qualitative results on some applications, including part segmentation, spatiotemporal tracking, and language-guided editing. We also provide ablation study results.
IV-A Experimental Setup
IV-A1 Datasets
@@
For comparisons on scene-level semantic segmentation, we select ScanNet [22] dataset as our 2D segmentation benchmark. ScanNet is a large-scale segmentation benchmark of indoor scenes, with calibrated RGBD trajectories, 3D point clouds, and ground-truth semantic label maps. We train 3D RGB Gaussians for all 1,201 scenes. We train our 3D semantic network on ScanNet train set and evaluate our performance on 12 scenes [45] from the validation set. To compare with closed-set methods, we follow the setting of [45] which maps ScanNet-20 classes to 21 classes from COCO dataset.
For 3D object localization tasks, we choose LERF [5] dataset as our benchmark. LERF dataset provide several 3D scenes containing long-tail objects and multi-scale semantics. The dataset is captured by iPhone App Polycam to get multi-view images and SfM points. In our experiments, we follow the setting of [12] to evaluate the localization accuracy on 4 different scenes.
@@
For qualitative evaluations, we choose MVImgNet [23] dataset as our part segmentation dataset, and CMU Panoptic [52] dataset as our spatiotemporal tracking dataset. MVImgNet is a multi-view single-object dataset containing 238 classes of objects with camera parameters and sparse point clouds. CMU Panoptic dataset is a large-scale dataset for multi-people engaging in social activities. For language-guided editing, we choose some scenes in the Mip-NeRF 360 [53] dataset to show our performance.
IV-A2 Implementation Details
All our experiments are trained on a NVIDIA RTX 4090 GPU. We train 10000 iterations for RGB Gaussians and 100 epochs for 3D semantic network. For scene-level semantic segmentation experiments, We apply LSeg to generate pixel-level open-vocabulary semantic features for 2D projection. As for 3D semantic network, we use MinkowskiNet34A [21] as our backbone. For 3D object localization tasks, we both apply CLIP with SAM and LSeg as our 2D pretrained model.
For part segmentation and spatiotemporal tracking, we use VLPart as our projection model. To evaluate our Semantic Gaussians on 4D Gaussians, we follow the work of Dynamic 3D Gaussians [24] to obtain dynamic Gaussians with temporal information.
IV-B Quantitative Results
@@
TABLE I: 2D semantic segmentation results on 12 scenes in ScanNet validation set. We report the mean IoU and the mean accuracy on all classes.
      Method Backbone mIoU mAcc
      closed-set methods
      Mask2Former [54] ViT 46.7 -
      DM-NeRF [55] NeRF 49.5 -
      SemanticNeRF [26] NeRF 59.2 -
      Panoptic Lifting [45] NeRF 65.2 -
      open-vocabulary methods
      OpenSeg [19] EfficientNet 53.4 75.1
      LSeg [18] ViT 56.1 74.5
      LERF [5] NeRF+CLIP 31.2 61.7
      PVLFF[56] NeRF+LSeg 52.9 67.0
      LangSplat [12] 3DGS+CLIP 24.7 42.0
      Feature3DGS [13] 3DGS+LSeg 59.2 75.1
      Ours 2D 3DGS+LSeg 61.0 76.6
      Ours 3D 3DGS+LSeg 59.7 74.7
      Ours 2D+3D 3DGS+LSeg 62.0 77.0
TABLE II: 3D object localization results on LERF dataset. We follow LangSplat [12] to report localization accuracy (%) on 4 scenes.
Scene LSeg LERF LangSplat Ours (LSeg) Ours (CLIP)
ramen 14.1 62.0 73.2 21.1 76.8
figurines 8.9 75.0 80.4 10.7 83.1
teatime 33.9 84.8 88.1 32.2 89.8
waldo_kitchen 27.3 72.7 95.5 31.8 90.9
overall 21.1 73.6 84.3 24.0 85.2
TABLE III: Performance of ablation studies.
       Setting mIoU mAcc
       original 62.0 77.0
       XYZ+RGB features 58.9 74.7
       20% Gaussian points 61.2 76.1
       10% input views 59.8 75.2
IV-B1 Open-Vocabulary Semantic Segmentation
@@
(a) basketball
(b) juggle
(c) softball
(d) tennis
Figure 5: Qualitative results of spatiotemporal tracking on the CMU Panoptic dataset. We choose 4 scenes with humans and dynamic objects to show the tracking performance.
We first evaluate our approach on the scene-level semantic segmentation task. We compare our method with several methods, including closed-set and open-vocabulary segmentation methods. We choose both 2D segmentation models and 3D segmentation methods based on NeRF or 3DGS. We report mIoU and mAcc as the metrics of semantic segmentation on the ScanNet-20 benchmark. We compare our method on ScanNet dataset. ScanNet dataset has several canonical classes (wall, floor, table, etc.) with ground truth semantic labels on each posed 2D image. As we only process off-the-shelf 3D Gaussians and do not care about their training process, we do not need to report metrics about rendering quality. The training of 3D Gaussian Splatting follows the official setting. In our method, we extract semantic features from LSeg to conduct 2D projection and 3D network training. Consequently, the result in this section will show how much our method will improve from 2D vision-language models.
Table I shows the performance of different methods. It can be observed that our Semantic Gaussians surpasses all open-vocabulary methods in mIoU and mAcc, and closely approaches the performance of state-of-the-art closed-set methods. This demonstrates that our approach effectively facilitates 3D scene understanding. We can also observe that both our 2D projection (Sec. III-B) and our 3D network (Sec. III-C) surpass the pre-trained LSeg model, even though all of our knowledge comes from it and relies on no ground truth labels. We assume the reason lies in the multi-view information integration, keeping the prediction consistent and mitigating some errors in low-quality views. Moreover, we notice that though the mIou and mAcc of our 3D network are lower than the 2D projection, the 2D and 3D ensemble will further improve our performance. We conjecture that some objects in certain scenes cannot be correctly recognized by 2D models in all views due to their low quality, while the 3D network could recognize them by utilizing geometric details.
Fig. 3 shows the segmentation results of open-vocabulary methods based on NeRF and 3DGS on the ScanNet dataset. As shown, LERF and LangSplat exhibit low segmentation accuracy. This is primarily because they utilize multi-scale CLIP features for scene understanding, while CLIP features struggle to align precisely with the scene at the pixel level, making them unsuitable for generating pixel-level semantic segmentation maps. On the other hand, PVLFF, Feature 3DGS, and Semantic Gaussians all employ LSeg to achieve open-vocabulary scene understanding. PVLFF performs relatively poorly and also suffers from slow rendering speeds as a NeRF-based method. Feature 3DGS and our method yield very similar performance, with Feature 3DGS even achieving more precise segmentation in certain views. However, Feature 3DGS requires retraining the semantic 3DGS for each scene, which is less efficient and flexible compared to our method.
IV-B2 Object Localization
Subsequently, we evaluated our method on the 3D object localization task across 4 scenes. Note that the LSeg model performed poorly on this task, significantly lagging behind CLIP-based methods like LERF and LangSplat. This likely lies in the fact that LSeg lacks the ability to recognize and segment long-tail objects, making it ineffective in locating objects within this dataset. Therefore, results of other LSeg-based methods are not included in the table. Fortunately, our method is not restricted by the 2D pre-trained model and can utilize SAM as a mask generator, combining it with the CLIP model to achieve versatile projection. Our Semantic Gaussians based on SAM+CLIP outperformed in 3 out of 4 scenarios in the segmentation task and achieved the highest average accuracy, demonstrating the effectiveness and flexibility of our method.
Figure 6: Visualization performance of VLPart [20] on CMU Panoptic dataset. The failure cases are highlighted by red boxes.
IV-B3 Ablation Study
We conduct ablation studies to figure out the efficiency of our method. As the performance of 2D projection and 3D network is presented in Sec. IV-B1, in this section, we ablate the performance of reducing 3D network input features, Gaussian points and input views. Points in 3D Gaussian Splatting have many features, i.e., coordinates, colors, rotations, scales, and opacities. If we only use coordinates and colors, they degrade to point clouds. Table III shows the comparisons of different inputs. We observe that if we reduce the input features, the mIoU and mAcc of our 3D networks will become much lower. This implies that the extra features in 3D Gaussian Splatting are important, and provide more information than RGB point clouds. Table III also shows the impact of reducing the number of Gaussian points and the number of input views, although these effects are less significant than reducing the input dimensions of the 3D network. This indicates that our method maintains a certain level of robustness even when there are fewer Gaussian points or input views, without suffering severe performance degradation.
IV-C Qualitative Evaluations
IV-C1 Part Segmentation
@@
IV-C2 Spatiotemporal Tracking
@@
Fig. 5 shows the qualitative result of our spatiotemporal tracking. As the scene often contains 12̃ humans and a dynamic object, we use VLPart in 2D projection and perform human part segmentation simultaneously. For each dynamic 3D Gaussian, we conduct our method at every frame to avoid looking at future frames. We render some novel views to show our capability of spatial tracking. The results show that our Semantic Gaussians can track human parts and objects with a high accuracy between different views and timesteps.
We also tried to segment each image individually from the same viewpoint by VLPart, and the results are shown in the Fig. 6. It can be observed that, although VLPart is able to produce fine boundaries, it makes different errors at different viewpoints or timesteps (highlighted in red boxes in the figure). This lack of spatiotemporal consistency in VLPart’s results makes it difficult to achieve spatiotemporal tracking. In contrast, our method, which is based on 4D Gaussians as the representation, inherently possesses strong consistency, enabling more accurate tracking.
Figure 7: Visualization results of instance segmentation results on 3 different scenes. We show the segmentation result of DEVA and our Semantic Gaussians. Different colors in the segmentation map denote different instances.
@@
IV-C3 Scene-Level Instance Segmentation
In this section, we will show the performance of Semantic Gaussians on scene-level instance segmentation tasks. Some related works [58, 59, 60] utilize SAM to generate weakly supervised masks and endow 3DGS with instance segmentation capabilities by redesigning the loss function and training 3DGS. Similarly, our method can leverage the knowledge from SAM to assign instance labels to 3DGS through a single projection. Specifically, following the approach of [58], we use DEVA [61] as the scene-level mask generator. DEVA is capable of assigning a unique scene-level ID to each object in a video, and we project this ID in a one-hot manner directly onto 3DGS. By rendering the semantic channels of 3DGS, instance segmentation maps from any viewpoint can be obtained.
Fig. 7 illustrates the visualization results of our method on the instance segmentation task. We choose 3 scenes from different datasets including LERF [5], 3D-OVS [4] and Mip-NeRF 360 [53]. It can be observed that our method can accurately delineate the boundaries of foreground objects with high precision without any retraining of 3DGS. However, for background objects or those that appear infrequently, our method fails to yield accurate boundaries and instance classifications. The primary reason is that DEVA, as a pre-trained model, cannot ensure the consistency of IDs for the background across different viewpoints. Our method can combine the segmentation results of DEVA from different viewpoints to achieve more consistent 3D instance segmentation results.
IV-C4 Language-guided Editing
In this section, we show the application of language-guided editing of our Semantic Gaussians. There are several works [58, 62, 63, 64, 65, 13] that segments 3DGS by SAM and conduct instance editing, but they often require the retraining of 3DGS. Our method can predict semantic embeddings for each 3D Gaussian point, and thus we can choose certain points by language query. We define some canonical operations such as removing, moving and color changing, and employ a language encoder to select the target Gaussians.
Fig. 8 shows some qualitative examples of language-guided editing on the room scene in the Mip-NeRF 360 dataset [53]. In these examples, we use CLIP text embeddings of ”glass bottle”, ”metal bowl” and ”slippers” to query and choose certain 3D Gaussian points in the scene, and we edit them by modifying their properties such as coordinates, colors, opacities, etc. From these examples, we observe that the language guidance can accurately select the target object, and thus we can perform various editing operations on the selected 3D Gaussians.
V Conclusion
In this work, we propose Semantic Gaussians, a novel approach to open-vocabulary 3D scene understanding via 3D Gaussian Splatting. Semantic Gaussians distill knowledge from pre-trained 2D encoders by projecting 2D pixel-level embeddings to 3D Gaussian points. Moreover, we introduce a 3D sparse convolutional network to predict semantic components with the input of RGB Gaussians, thus achieving zero-shot generalization to unseen 3D scenes. We conduct experiments on the ScanNet segmentation benchmark to prove its effectiveness and exhibit downstream applications such as part segmentation, spatiotemporal tracking, instance segmentation, and scene editing. Our work paves the way for real-world applications of 3D Gaussian Splatting, such as embodied agents and augmented reality systems.